# NaviAgent: Bilevel Planning on Tool Dependency Graphs for Function Calling

### v1

@@ 9. Use call_api only when:
    - The target_api is not in the list of successfully executed APIs
@@ E.3.3 API Simulator Prompts
Act as an API Chain Simulator to generate responses based on historical call
chains.
@@ 3. Output Format (JSON):
  "status": "<success>", // Always ’success’ per operation completion
@@ E.3.4 Simulated User Response Agent Prompts
As an API chain parameter extractor, directly return exact parameter values
from the given API workflows without any modification.
@@ 1. Parameter Extraction Priority
    Always return raw parameter values from the latest API call
    Return empty string for blank parameters (e.g. param-cuisines_1 -> "")
@@ 2. Response Requirements
API Context: [param-trial_phase_1:’Phase 2’ param-calories_max_1:’2000’...]
@@
Response: I don’t Know international roaming activation information.
@@ E.4.1 Evaluation Prompts
As an expert in response quality evaluation, you need to perform the
following steps:
@@ 1. Reference Path Analysis
- Understand the simulated nature of reference API call paths.
- Be aware of potential discrepancies: API names/parameter formats may
  differ from actual implementations.
@@ VI. Input Data Interface
Appendix F Experiments on API-Bank
@@
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B Decider 56.1 38.6 4.37 68.2 31.2 4.94 76.3 34.6 5.20 70.5 33.8 4.98 D+N(Alpha) 64.9 49.1 4.83 72.7 36.4 5.16 68.7 36.0 5.03 69.8 37.8 5.05 D+N(Heur) 73.7 56.1 4.88 67.0 40.9 5.25 66.8 33.6 5.46 67.8 39.4 5.27 Qwen2.5-32B Decider 68.4 50.9 4.17 84.1 43.2 5.59 72.5 36.5 6.01 76.6 41.0 5.54 D+N(Alpha) 89.5 68.4 4.47 80.1 35.8 5.70 85.3 45.0 6.03 83.8 44.4 5.62 D+N(Heur) 91.2 73.7 4.55 87.5 46.0 5.73 84.8 45.5 5.63 86.7 49.3 5.46 Deepseek-R1-32B Decider 91.2 71.9 5.33 73.3 36.4 6.44 72.5 30.8 6.42 75.2 38.3 6.16 D+N(Alpha) 89.5 63.2 5.63 90.9 48.3 6.59 81.5 34.1 6.43 86.3 43.5 6.35 D+N(Heur) 99.7 77.3 5.03 89.6 46.8 6.21 79.2 35.2 6.80 86.0 45.2 6.17 Deepseek-V3 Decider 89.5 73.7 4.80 90.9 52.3 5.47 92.4 47.9 5.71 91.4 52.9 5.45 D+N(Alpha) 99.9 82.5 4.67 98.9 58.5 6.05 99.1 51.2 5.54 99.1 58.1 5.58 D+N(Heur) 99.9 89.5 5.10 99.4 61.4 5.74 99.1 52.6 5.72 99.3 60.8 5.61 GPT-4o Decider 99.9 80.7 4.28 99.4 56.4 4.96 98.6 46.5 4.93 99.1 54.8 4.82 D+N(Alpha) 99.9 78.5 4.04 98.9 56.4 4.98 98.6 52.2 4.57 98.9 57.2 4.64 D+N(Heur) 99.9 74.3 3.75 98.3 59.1 5.02 98.1 57.9 4.90 98.4 60.5 4.77 Qwen2.5-14B(SFT) Decider 80.7 52.6 4.57 80.7 41.6 5.12 79.6 45.5 5.04 80.2 44.6 5.00 D+N(Alpha) 94.7 64.3 4.57 89.8 44.1 5.16 88.2 42.2 5.01 89.6 45.5 4.99 D+N(Heur) 93.0 72.1 4.74 89.8 48.6 5.22 87.7 45.0 5.02 89.2 49.8 5.04
Table 5: Impact of Navigator and Search Algorithms on API-Bank. Metrics are reported as in Table 2.

### v2

@@ 9. Use call_api only when:
    - The target_api is not in the list of successfully
    executed APIs
@@ E.3.3 API Simulator Prompts
Act as an API Chain Simulator to generate responses based on
historical call chains.
@@ 3. Output Format (JSON):
  "status": "<success>", // Always ’success’ per operation
  completion
@@ E.3.4 Simulated User Response Agent Prompts
As an API chain parameter extractor, directly return exact
parameter values from the given API workflows without any
modification.
@@ 1. Parameter Extraction Priority
    Always return raw parameter values from the latest API
    call
    Return empty string for blank parameters (e.g. param-
    cuisines_1 -> "")
@@ 2. Response Requirements
API Context: [param-trial_phase_1:’Phase 2’ param-
calories_max_1:’2000’...]
@@
Response: I don’t Know international roaming activation
information.
@@ E.4.1 Evaluation Prompts
As an expert in response quality evaluation, you need to
perform the following steps:
@@ 1. Reference Path Analysis
- Understand the simulated nature of reference API call
paths.
- Be aware of potential discrepancies: API names/parameter
formats may differ from actual implementations.
@@ VI. Input Data Interface
Appendix F Experiments
F.1 Evaluating Our Approach on Real-World APIs
To further validate our framework, we conducted real-world evaluations on 50 APIs from RapidAPI, covering weather, air quality, restaurants, real estate, geolocation, hotels, and sports. A total of 60 queries (20 easy, 20 medium, and 20 hard) were carefully designed to ensure comprehensive coverage across these domains. On this real-world testbed, our framework consistently outperformed the $\alpha$-UMI baseline in both effectiveness and efficiency. Metrics are reported as in Table 5 and Time is measured in seconds (s).
Model Method TCR TSR Steps Time
Qwen2.5-14B ReAct 31.7 21.7 3.70 16
ToolLLM 53.3 23.3 4.10 19
$\alpha$-UMI 76.7 31.7 5.80 27
NaviAgent 65.0 36.7 4.90 25
Qwen2.5-32B ReAct 35.0 25.0 3.80 19
ToolLLM 48.3 30.0 4.00 23
$\alpha$-UMI 78.3 41.6 6.04 32
NaviAgent 86.7 53.3 4.84 27
DeepSeek-V3 ReAct 55.0 33.3 3.75 23
ToolLLM 53.3 35.0 4.05 25
$\alpha$-UMI 85.0 48.3 6.17 39
NaviAgent 98.3 63.3 5.11 35
Table 5: Real-World APIs Test.
F.2 Runtime Experiments
Table 6 presents the runtime (in seconds) of NaviAgent variants across different models on ToolBench. Notably, the Dynamic+A method consistently achieves lower runtime across all models, with the most significant improvement observed in Deepseek-V3: compared to the Base method (55.8 seconds), Dynamic+A reduces the runtime by 15 seconds, corresponding to a relative improvement of approximately 26.9%. Among all methodological variants, Dynamic+H demonstrates the optimal overall performance; however, it is constrained by higher runtime induced by heuristic strategies and excessive search scale, which will be the focus of subsequent optimization efforts.
Model Method Easy Medium Hard All
Qwen2.5-14B Base 26.6 34.0 44.5 34.0
Static+A 21.6 27.1 36.1 27.4
Dynamic+A 19.8 25.3 34.4 25.6
Dynamic+H 22.9 30.4 39.6 30.1
Qwen2.5-32B Base 33.4 41.7 53.3 41.7
Static+A 24.0 33.0 41.1 32.0
Dynamic+A 24.0 31.0 38.6 30.5
Dynamic+H 28.1 33.8 48.4 35.3
Deepseek-R1-32B Base 36.2 44.2 61.2 45.5
Static+A 27.5 33.7 46.8 34.7
Dynamic+A 24.6 34.6 44.5 33.6
Dynamic+H 31.0 36.7 49.7 37.8
Deepseek-V3 Base 43.6 56.6 71.3 55.8
Static+A 34.4 44.5 55.5 43.8
Dynamic+A 30.3 41.6 53.3 40.6
Dynamic+H 37.0 47.3 61.6 47.3
GPT-4o Base 42.3 55.6 75.6 55.9
Static+A 34.9 44.1 59.8 44.7
Dynamic+A 32.6 43.8 56.4 43.1
Dynamic+H 36.9 47.1 61.5 47.1
Qwen2.5-14B(SFT) Base 27.0 38.0 50.1 37.2
Static+A 22.3 27.1 37.8 28.0
Dynamic+A 19.9 27.9 35.9 27.2
Dynamic+H 24.5 31.4 40.6 31.3
Table 6: Runtime(in Seconds) of NaviAgent Variants on ToolBench
F.3 Experiments on API-Bank
Table 7 demonstrates that the experimental outcomes of the API-Bank dataset are consistent with those observed in the ToolBench-based experiments.
@@ F.3 Experiments on API-Bank
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B Base 47.8 33.4 5.40 60.5 24.8 6.09 71.6 29.9 6.47 63.1 27.6 6.06 Static+A 63.4 44.8 4.88 72.4 32.8 5.38 68.4 34.3 5.41 67.9 34.0 5.22 Dynamic+A 64.9 49.1 4.93 72.7 36.4 5.32 68.7 36.0 5.36 68.3 36.7 5.18 Dynamic+H 73.1 56.1 4.71 66.6 40.4 5.27 66.4 33.5 5.63 65.7 37.9 5.26 Qwen2.5-32B Base 61.6 46.6 5.26 78.6 35.0 6.84 68.9 30.7 7.38 70.4 33.4 6.78 Static+A 88.6 65.7 4.63 80.0 34.3 5.90 84.6 39.9 6.29 81.3 39.5 5.82 Dynamic+A 89.5 68.4 4.57 80.1 35.8 5.86 85.3 45.0 6.36 81.8 42.8 5.83 Dynamic+H 90.8 74.0 4.54 87.0 44.7 5.70 84.2 45.5 5.66 84.1 47.2 5.43 Deepseek-R1-32B Base 88.2 66.2 6.41 65.3 28.3 7.79 64.5 25.4 7.83 65.9 30.3 7.49 Static+A 89.2 60.7 5.77 88.1 46.1 6.83 81.2 30.5 6.82 83.0 39.2 6.56 Dynamic+A 89.5 63.2 5.73 90.9 48.3 6.75 81.5 34.1 6.76 84.2 42.0 6.49 Dynamic+H 99.1 77.6 4.96 89.4 46.9 6.06 79.3 34.4 6.81 83.6 43.2 6.16 Deepseek-V3 Base 86.3 67.0 5.95 86.3 46.5 6.65 85.4 42.1 7.09 83.9 45.5 6.64 Static+A 97.7 77.4 4.84 98.6 55.5 6.17 98.8 48.1 5.82 96.4 53.1 5.72 Dynamic+A 99.9 82.5 4.77 98.9 58.5 6.21 99.1 51.2 5.87 96.9 56.3 5.76 Dynamic+H 98.8 88.9 5.00 98.0 60.0 5.74 98.6 52.3 5.88 96.2 58.0 5.60 GPT-4o Base 96.8 74.9 5.23 92.4 48.3 6.15 94.5 38.6 6.19 91.8 45.4 5.93 Static+A 99.6 76.8 4.15 98.6 54.5 5.17 98.3 46.9 4.85 96.3 52.0 4.79 Dynamic+A 99.9 78.5 4.14 98.9 56.4 5.14 98.6 52.2 4.90 96.6 55.5 4.80 Dynamic+H 98.9 76.1 3.70 98.1 57.9 5.00 97.0 57.8 5.00 95.5 58.5 4.75 Qwen2.5-14B(SFT) Base 76.0 45.4 5.63 74.9 35.1 6.35 76.3 40.6 6.39 74.0 38.0 6.15 Static+A 94.1 60.6 4.69 88.7 41.3 5.24 87.9 41.3 5.28 86.9 42.4 5.08 Dynamic+A 94.7 64.3 4.67 89.8 44.1 5.32 88.2 42.2 5.34 87.5 44.3 5.14 Dynamic+H 93.2 71.0 4.61 90.2 48.3 5.17 87.6 44.5 5.14 87.3 47.8 4.98
Table 7: Impact of NaviAgent Variants on API-Bank. Metrics are reported as in Table 2.